An uncharged molecule of DNA (deoxyribonucleic acid) is 2.17 µm long. The ends of the molecule become singly ionized so that there is a charge of −1.6 × 10−19 C on one end and +1.6 × 10−19 C on the other. The helical molecule acts like a spring and compresses 1.3% upon becoming charged. Find the effective spring constant of the molecule. The value of Coulomb’s constant is 8.98755 × 109 N · m2 /C 2 and the acceleration due to gravity is 9.8 m/s 2 . Answer in units of N/m.
Answer #1

ginen d=2-17x100 91= 4.641019 02=+1.68109 al = 1.34. thers then [f=ka = k1192 KU - k( 0.013 (249x106) = (8.987551109) (1.6X1